# Similarity-Distance-Magnitude Activations

### v1

@@ Abstract
We introduce a more robust and interpretable formulation of the standard softmax activation function commonly used with neural networks by adding Similarity (i.e., correctly predicted depth-matches into training) awareness and Distance-to-training-distribution awareness to the existing output Magnitude (i.e., decision-boundary) awareness. When used as the final-layer activation with language models, the resulting Similarity-Distance-Magnitude (sdm) activation function is more robust than the softmax function to co-variate shifts and out-of-distribution inputs in high-probability regions, and provides interpretability-by-exemplar via dense matching. Complementing the prediction-conditional estimates, the sdm activation enables a partitioning of the class-wise empirical CDFs to guard against low class-wise recall among selective classifications. These properties make it preferable for selective classification, even when considering post-hoc calibration methods over the softmax.
@@ 1 Introduction
Large neural-network-based language models (LLMs) pose a challenge for interpretable and reliable deployment given the non-identifiability of their parameters (Hwang & Ding, 1997, inter alia) [^1] Informally, this means that two or more distinct sets of values for the parameters can result in identical output distributions. As a consequence, interpreting the parameters of such models is typically much more complicated than with a simple linear regression model, for example., which can number in the billions or more. Instead of directly interpreting parameters, we might instead consider moving the focus of interpretation to auditing predictions as a form of interpretability by example, or exemplar, over the representation space of such models via dense matching. However, for real-world deployments, robust approaches for predictive uncertainty are also needed, both for human decision-making and for constructing sequentially dependent LLM pipelines.
@@
Known theoretical results limit the statistical quantities that can be derived over LLMs. Statistical assurances in the distribution-free setting are limited to approximately conditional quantities (Valiant, 1984; Lei & Wasserman, 2014; Foygel Barber et al., 2020, inter alia). Further, even typical approximately conditional quantities can be difficult to obtain in practice, since the minimal assumption of exchangeability with a known held-out data set is itself often violated with co-variate and label shifts, which can be difficult to foresee with existing methods. Epistemologically, the prevalence of hallucinations and highly-confident wrong answers with widely deployed LLMs suggests a technical impasse in effectively modeling the predictive uncertainty, despite significant work from Bayesian, Frequentist, and empirically motivated perspectives (Gal & Ghahramani, 2016; Angelopoulos et al., 2021; Guo et al., 2017; Lakshminarayanan et al., 2017; Ovadia et al., 2019, inter alia). A foundational piece is evidently missing from the picture.
@@
Given these intrinsic challenges, we approach the problem of uncertainty quantification over LLMs from a new angle and ask: Can we leverage the metric learning and dense matching capabilities of neural networks over high-dimensional inputs to at least aim to maximize, with minimal distributional assumptions, the separation of aleatoric (irreducible) uncertainty and epistemic (reducible) uncertainty, decomposing the sources of the latter in a manner that is interpretable and actionable?
@@
We answer this question in the affirmative with a conceptually parsimonious, LLM-driven partitioning of the data to decompose sources of epistemic uncertainty: Correctly predicted depth-matches into the training set ($\operatorname{\text{{Similarity}}}$), the $\operatorname{\text{{Distance}}}$ to the training set, and the distance to the decision-boundary ($\operatorname{\text{{Magnitude}}}$). We use these signals to construct a new activation function, the $\operatorname{\text{{sdm}}}$ activation, which can be used as a replacement for the standard $\operatorname{softmax}$ operation, as, for example, the final-layer activation. The $\operatorname{\text{{sdm}}}$ activation enables more reliable estimates of the predictive uncertainty for selective classification (Chow, 1957; Geifman & El-Yaniv, 2017, inter alia), which addresses the need for uncertainty quantification with LLMs as part of multi-stage decision pipelines, in settings subject to co-variate shifts and out-of-distribution inputs.
@@
  • We introduce the $\operatorname{\text{{Similarity}}}$-$\operatorname{\text{{Distance}}}$-$\operatorname{\text{{Magnitude}}}$ ($\operatorname{\text{{sdm}}}$) activation function, which encodes strong signals of epistemic uncertainty, to replace the $\operatorname{softmax}$ operation.
  • We specifically examine the behavior of the $\operatorname{\text{{sdm}}}$ activation as a final-layer activation over language models for selective classification. Even when considering typical post-hoc calibration methods, the $\operatorname{\text{{sdm}}}$ activation is more robust than the $\operatorname{softmax}$ activation to co-variate-shifted and out-of-distribution inputs, ceteris paribus, while remaining informative over in-distribution data.
  • We introduce an approach for controlling the recall of selective classifications, based on a data-driven partitioning of the class-wise CDFs via the $\operatorname{\text{{sdm}}}$ activation. Combined with the empirically-robust prediction-conditional estimates from the $\operatorname{\text{{sdm}}}$ activation, this results in a straightforward, and empirically achievable, quantity of interest for selective classification.
@@ 2.1 Setting
We consider the standard multi-class classification setting (e.g., predicting the binary sentiment of a movie review). We are given a training dataset, $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}=\{({\bm{x}}_{n},y_{n})\}_{n=1}^{N}$ of inputs, ${\bm{x}}\in{\mathcal{X}}$, paired with their corresponding ground-truth discrete labels, $y\in{\mathcal{Y}}=\{1,\ldots,C\}$, and a labeled calibration dataset, $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$, drawn from the same distribution as $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$. We are then given a new test instance, $\bm{x}$}, from an unlabeled test set, $\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}$, and seek to estimate the label with a prediction, $\hat{y}$, via the un-normalized log probabilities (“logits”, informally) of a final linear layer: ${\bm{z}}={\bm{W}}^{T}{\bm{h}}+{\bm{b}}$, where ${\bm{h}}=\operatorname{{\rm{network}}}({\bm{x}};\theta)$ is the final hidden state of a network parameterized by $\theta$. The discrete prediction is taken as $\hat{y}=\operatornamewithlimits{arg\,max}{{\bm{z}}}$; however, for learning $\theta$, $\bm{W}$}, and $\bm{b}$}, and for human decision-making, we also seek an estimate of the predictive uncertainty, $p(y\,|\,{\bm{x}})$, which is typically obtained by normalizing $\bm{z}$} via the $\operatorname{softmax}$ operation described next. We will make a distinction between models, $\mathcal{M}$} (defined by $\theta$, $\bm{W}$}, and $\bm{b}$}, and when applicable, the exemplar adaptor, described below), which produce the prediction, $\hat{y}$, and estimators, $\mathcal{E}$}, which provide an estimate of $p(y\,|\,{\bm{x}})$, because different estimators can be used over the same model.
@@ 2.2 Softmax and the Cross-Entropy loss
The $\operatorname{softmax}$ activation is commonly used in neural network architectures, including, for example, as a router in self-attention mechanisms (Vaswani et al., 2017) and mixture-of-experts models (Shazeer et al., 2017), and forming the basis of the cross-entropy loss used for next-token training of LLMs. It is the typical final output layer of LLMs, converting the un-normalized model logits to a normalized probability distribution:
@@ 3 Methods
with a corresponding negative log likelihood loss that takes into account the change of base (§ 3.1). Unique to this setting, a modification to label-conditional conformal prediction (Vovk et al., 2005) then follows via a parsimonious partitioning of the class-wise empirical CDFs, providing a principled basis for controlling the recall among selective classifications, combined with empirically-robust prediction-conditional estimates.
@@ 3.1.1 Exemplar Adaptor
We take as the CNN of our exemplar adaptor $g:{\bm{h}}\in\mathbb{R}^{D}\mapsto{\bm{h}}^{\prime}\in\mathbb{R}^{M}$, a 1-D CNN that takes as input $\bm{h}$} of the underlying network. The CNN has $M$ filters, the filter applications of which produce $\bm{h}$$^{\prime}$, the distilled representation of the underlying network. A final linear layer, ${\bm{z}}^{\prime}={\bm{W}}^{\prime T}{\bm{h}}^{\prime}+{\bm{b}}^{\prime},{\bm{z}}^{\prime}\in\mathbb{R}^{C}$, then replaces the underlying network’s linear layer, with the discrete prediction taken as $\hat{y}=\operatornamewithlimits{arg\,max}{{\bm{z}}^{\prime}}$. This exemplar adaptor will then enable us to derive the key signals of epistemic uncertainty, $\operatorname{\text{{Similarity}}}$, $\operatorname{\text{{Distance}}}$, and $\operatorname{\text{{Magnitude}}}$ described next.
@@ 3.1.5 SDM Activation: Formulation
The output distribution becomes sharper with higher values of $\operatorname{\text{$q$}}$, $\operatorname{\text{$d$}}$, and $z^{\prime}$. Also note that when $\operatorname{\text{$d_{\rm{nearest}}$}}$ exceeds the largest distance observed in the labeled data, $\operatorname{\text{$d$}}=0$ and the output distribution is uniform, reflecting a maximally high (i.e., out-of-distribution) epistemic uncertainty estimate. The standard $\operatorname{softmax}$ with $\tau=1$ is recovered by setting q=e2,d=1q=e-2,d=1. As with the $\operatorname{softmax}$ operation, $\operatornamewithlimits{arg\,max}\operatorname{\text{{sdm}}}({\bm{z}}^{\prime})=\operatornamewithlimits{arg\,max}{\bm{z}}^{\prime}$.
@@ 3.1.6 SDM Activation: Loss and Training
The first epoch of training is initialized with a standard $\operatorname{softmax}$ (i.e., setting q=e2,d=1q=e-2,d=1). Training then proceeds by re-calculating $q$ and $d$ for each ${\bm{x}}\in\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ after each epoch. We take as the stopping criteria for one learning round as the epoch with the highest balanced (across classes) average loss over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$. We repeat this process for $J$ iterations of random shuffles and splits of $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ and $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ and parameter initializations, choosing the final model as that with the globally highest balanced (across classes) average loss over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$.
@@ 3.2 Evaluating Selective Classification
As a common, unambiguous baseline quantity for comparing selective classifiers, we seek a straightforward balance between precision and recall. Among the selective classifications from an estimator, we seek (I) prediction-conditional accuracy and (II) class-conditional accuracy above a given threshold, $\alpha\in(\frac{1}{C},1]$.
To evaluate this quantity, we only consider the points for which the given estimator assigns a high-probability of at least $\alpha$, which is typically near 1, such as $\alpha=0.95$ in our experiments. We refer to this set of points as the admitted, or non-rejected, set. Then, given ground-truth values for $\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}$, we assess whether the conditional accuracies of the admitted set are at least $\alpha$ when (I) stratifying by the predicted labels, $\hat{y}$, and when (II) stratifying by the true labels, $y$.
@@
Quantity I corresponds to top-label calibration (Gupta & Ramdas, 2022), but with a single bin for evaluation, $\alpha,1]$, removing ambiguity with regard to the choice of binning the probabilities. Quantity II does not directly correspond to quantities typically examined in the empirical calibration literature (Brier, 1950; Dawid, 1982; Guo et al., 2017; Vaicenavicius et al., 2019; Kull et al., 2019, inter alia), but it approximates444 For simplicity of presentation, we omit consideration of the Beta-distributed error term that is a function of the effective sample size of split-conformal coverage (Vovk, 2012). In practice, this term is negligible in the present experiments given $|\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}|$ and the resolution of the comparisons. label-conditional conformal coverage in the restricted setting of class-wise thresholds that only admit prediction sets of cardinality 1. We describe a straightforward procedure to estimate this quantity next.
3.2.1 Controlling Selective Classification Recall
In general, the statistical coverage assurance of marginal split-conformal estimators does not directly correspond to selection, since the coverage assurance is not conditional on the set size. We may instead seek one of various approximately conditional notions of coverage (Romano et al., 2020; Angelopoulos et al., 2021, inter alia); however, there is no guarantee that when we stratify by sets of cardinality 1, coverage will be maintained. However, there is a restricted setting in which label-conditional conformal estimators do provide a meaningful notion of class-conditional coverage for selection. Assuming $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ and $\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}$ are exchangeable, if the conformity score for each label is from a categorical distribution and the resulting thresholding of the class-wise empirical CDFs results in class-wise thresholds that are all greater than $\frac{1}{C}$, then the cardinality 1 sets will, on average, obtain class-conditional coverage, by definition. [^5] Although the formal interpretations are not identical, the evaluation of class-conditional coverage for a single estimate of such cardinality 1 prediction sets in this restricted setting over $\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}$ is numerically equivalent to assessing the class-conditional accuracy, when not considering the sample size error term. Unfortunately, it may be rare to encounter this restricted setting over the full data distributions of real-world tasks. Instead, we will use the $\operatorname{\text{{sdm}}}$ activation to estimate a partitioning of the distribution into a region that approximately fulfills these assumptions.

### v2

@@ Abstract
We introduce the Similarity-Distance-Magnitude (sdm) activation function, a more robust and interpretable formulation of the standard $\operatorname{softmax}$ activation function, adding Similarity (i.e., correctly predicted depth-matches into training) awareness and Distance-to-training-distribution awareness to the existing output Magnitude (i.e., decision-boundary) awareness, and enabling interpretability-by-exemplar via dense matching. We further introduce the sdm estimator, based on a data-driven partitioning of the class-wise empirical CDFs via the $\operatorname{\text{{sdm}}}$ activation, to control the class- and prediction-conditional accuracy among selective classifications. When used as the final-layer activation over pre-trained language models for selective classification, the $\operatorname{\text{{sdm}}}$ estimator is more robust to co-variate shifts and out-of-distribution inputs than existing calibration methods using $\operatorname{softmax}$ activations, while remaining informative over in-distribution data.
@@ 1 Introduction
Neural-network-based language models (LMs) pose a challenge for interpretable and reliable deployment given the non-identifiability of their parameters (Hwang & Ding, 1997, inter alia) [^1] Informally, this means that two or more distinct sets of values for the parameters can result in identical output distributions. As a consequence, interpreting the parameters of such models is typically much more complicated than with a simple linear regression model, for example., which can number in the billions or more. Instead of directly interpreting parameters, one option is to move the focus of interpretation to auditing predictions as a form of interpretability by example, or exemplar, over the representation space of such models via dense matching. However, for real-world deployments, robust approaches for predictive uncertainty are also needed, both for human decision-making and for constructing sequentially dependent LM pipelines.
@@
Known theoretical results limit the statistical quantities that can be derived over LMs. Statistical guarantees in the distribution-free setting are limited to approximately conditional quantities (Valiant, 1984; Lei & Wasserman, 2014; Foygel Barber et al., 2020, inter alia). Further, even typical approximately conditional quantities can be difficult to obtain in practice, since the minimal assumption of exchangeability with a known held-out data set is itself often violated with co-variate and label shifts, which can be difficult to foresee with existing methods. Epistemologically, the prevalence of hallucinations and highly-confident wrong answers with widely deployed LMs suggests a technical impasse in effectively modeling the predictive uncertainty, despite significant work from Bayesian, Frequentist, and empirically motivated perspectives (Gal & Ghahramani, 2016; Angelopoulos et al., 2021; Guo et al., 2017; Lakshminarayanan et al., 2017; Ovadia et al., 2019, inter alia). A foundational piece is evidently missing from the picture.
@@
Given these intrinsic challenges, we approach the problem of uncertainty quantification over LMs from a new angle and ask: Can we leverage the metric learning and dense matching capabilities of neural networks over high-dimensional inputs to at least aim to maximize, with minimal distributional assumptions, the separation of aleatoric (irreducible) uncertainty and epistemic (reducible) uncertainty, decomposing the sources of the latter in a manner that is interpretable and actionable?
@@
We answer this question in the affirmative with a conceptually parsimonious, data-driven partitioning of the data to decompose sources of epistemic uncertainty: Correctly predicted depth-matches into the training set ($\operatorname{\text{{Similarity}}}$), the $\operatorname{\text{{Distance}}}$ to the training set, and the distance to the decision-boundary ($\operatorname{\text{{Magnitude}}}$). We use these signals to construct a new activation function, the $\operatorname{\text{{sdm}}}$ activation, which can be used as a replacement for the standard $\operatorname{softmax}$ activation, as, for example, the final-layer activation. The $\operatorname{\text{{sdm}}}$ activation enables more reliable estimates of the predictive uncertainty for selective classification (Chow, 1957; Geifman & El-Yaniv, 2017, inter alia), which addresses the need for uncertainty quantification with LMs used in multi-stage decision pipelines, in settings subject to co-variate shifts and out-of-distribution inputs.
@@
  • We introduce the $\operatorname{\text{{Similarity}}}$-$\operatorname{\text{{Distance}}}$-$\operatorname{\text{{Magnitude}}}$ ($\operatorname{\text{{sdm}}}$) activation function, which encodes strong signals of epistemic uncertainty, for use with neural networks.
  • We introduce the $\operatorname{\text{{sdm}}}$ estimator for use in controlling class- and prediction-conditional accuracy among selective classifications, based on a data-driven partitioning of the class-wise empirical cumulative distribution functions (eCDFs) over the output via the $\operatorname{\text{{sdm}}}$ activation.
  • We examine the behavior of the $\operatorname{\text{{sdm}}}$ activation as a final-layer activation over pre-trained language models, using the $\operatorname{\text{{sdm}}}$ estimator for selective classification. We demonstrate empirically that the $\operatorname{\text{{sdm}}}$ estimator is more robust to co-variate-shifts and out-of-distribution inputs than existing classes of post-hoc calibration methods over $\operatorname{softmax}$ activations, while remaining informative over in-distribution data.
@@ 2.1 Setting
We consider the standard multi-class classification setting (e.g., predicting the sentiment of a movie review). We are given a training dataset, $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}=\{({\bm{x}}_{n},y_{n})\}_{n=1}^{N}$ of inputs, ${\bm{x}}\in{\mathcal{X}}$, paired with their corresponding ground-truth discrete labels, $y\in{\mathcal{Y}}=\{1,\ldots,C\}$, and a labeled calibration dataset, $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$, drawn from the same distribution as $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$. We are then given a new test instance, $\bm{x}$}, from an unlabeled test set, $\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}$, and seek to estimate the label with a prediction, $\hat{y}$, via the un-normalized log probabilities (“logits”, informally) of a final linear layer: ${\bm{z}}={\bm{W}}^{T}{\bm{h}}+{\bm{b}}$, where ${\bm{h}}=\operatorname{{\rm{network}}}({\bm{x}};\theta)$ is the final hidden state of a network parameterized by $\theta$. The discrete prediction is taken as $\hat{y}=\operatornamewithlimits{arg\,max}{{\bm{z}}}$; however, for learning $\theta$, $\bm{W}$}, and $\bm{b}$}, and for human decision-making, we also seek an estimate of the predictive uncertainty, $p(y\,|\,{\bm{x}})$, which is typically obtained by normalizing $\bm{z}$} via the $\operatorname{softmax}$ activation described next. We will make a distinction between models, $\mathcal{M}$} (defined by $\theta$, $\bm{W}$}, and $\bm{b}$}, and when applicable, the exemplar adaptor, described below), which produce the prediction, $\hat{y}$, and estimators, $\mathcal{E}$}, which provide an estimate of $p(y\,|\,{\bm{x}})$, because different estimators can be used over the same model.
@@ 2.2 Softmax and the Cross-Entropy loss
The $\operatorname{softmax}$ activation is commonly used in neural network architectures, including, for example, as a router in self-attention mechanisms (Vaswani et al., 2017) and mixture-of-experts models (Shazeer et al., 2017), and forming the basis of the cross-entropy loss used for next-token training of LMs. It is the typical final output layer of LMs, converting the un-normalized model logits to a normalized probability distribution:
@@ 3 Methods
with a corresponding negative log likelihood loss that takes into account the change of base (§ 3.1). Unique to this setting, a modification to label-conditional conformal prediction (Vovk et al., 2005) then follows via a parsimonious partitioning of the class-wise empirical CDFs, providing a principled basis for controlling the class-conditional accuracy among selective classifications, combined with empirically-robust prediction-conditional estimates.
@@ 3.1.1 Exemplar Adaptor
We take as the CNN of our exemplar adaptor $g:{\bm{h}}\in\mathbb{R}^{D}\mapsto{\bm{h}}^{\prime}\in\mathbb{R}^{M}$, a 1-D CNN that takes as input $\bm{h}$} of the underlying network. The CNN has $M$ filters, the filter applications of which produce $\bm{h}$$^{\prime}$, the distilled representation of the underlying network. A final linear layer, ${\bm{z}}^{\prime}={\bm{W}}^{\prime T}{\bm{h}}^{\prime}+{\bm{b}}^{\prime},{\bm{z}}^{\prime}\in\mathbb{R}^{C}$, then replaces the underlying network’s linear layer, with the discrete prediction taken as $\hat{y}=\operatornamewithlimits{arg\,max}{{\bm{z}}^{\prime}}$. This exemplar adaptor will then enable us to derive the $\operatorname{\text{{Similarity}}}$, $\operatorname{\text{{Distance}}}$, and $\operatorname{\text{{Magnitude}}}$ values, as defined next.
@@ 3.1.5 SDM Activation: Formulation
The output distribution becomes sharper with higher values of $\operatorname{\text{$q$}}$, $\operatorname{\text{$d$}}$, and $z^{\prime}$. When $\operatorname{\text{$d_{\rm{nearest}}$}}$ exceeds the largest distance observed in the labeled data, $\operatorname{\text{$d$}}=0$ and the output distribution is uniform, reflecting maximally high uncertainty. The standard $\operatorname{softmax}$ with $\tau=1$ is recovered by setting q=e2,d=1q=e-2,d=1. As with the $\operatorname{softmax}$ activation, $\operatornamewithlimits{arg\,max}\operatorname{\text{{sdm}}}({\bm{z}}^{\prime})=\operatornamewithlimits{arg\,max}{\bm{z}}^{\prime}$.
@@ 3.1.6 SDM Activation: Loss and Training
The first epoch of training is initialized with a standard $\operatorname{softmax}$ (i.e., setting q=e2,d=1q=e-2,d=1). Training then proceeds by re-calculating $q$ and $d$ for each ${\bm{x}}\in\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ after each epoch. We take as the stopping criteria for one learning round as the epoch with the lowest balanced (across classes) average loss over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$. We repeat this process for $J$ iterations of random shuffles and splits of $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ and $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ and parameter initializations, choosing the final model as that with the globally lowest balanced (across classes) average loss over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$.
@@ 3.2 Evaluating Selective Classification
As a common, unambiguous baseline quantity for comparing selective classifiers over a held-out test set, $\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}$, we seek an easy-to-interpret and easy-to-evaluate metric, reflecting real-world applications. Among the selective classifications from an estimator, we seek (Quantity I) prediction-conditional accuracy at or above a given threshold, $\alpha\in(\frac{1}{C},1]$, and (Quantity II) class-conditional accuracy at or above that same threshold, $\alpha$.
To evaluate this metric, we only consider the points for which the given estimator assigns a high-probability of at least $\alpha$, which is typically near 1, such as $\alpha=0.95$ in our experiments. We refer to this set of points as the admitted, or non-rejected, set. Then, given ground-truth values for $\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}$, we assess whether the conditional accuracies of the admitted set are at least $\alpha$ when (Quantity I) stratifying by the predicted labels, $\hat{y}$, and when (Quantity II) stratifying by the true labels, $y$.
@@
Quantity I corresponds to top-label calibration (Gupta & Ramdas, 2022), but with a single bin for evaluation, $\alpha,1]$, removing ambiguity with regard to the choice of binning the probabilities. Quantity II does not directly correspond to quantities typically examined in the calibration literature (Brier, 1950; Dawid, 1982; Guo et al., 2017; Vaicenavicius et al., 2019; Kull et al., 2019, inter alia), but it approximates [^4] For simplicity of presentation, we omit consideration of the Beta-distributed error term that is a function of the effective sample size of split-conformal coverage (Vovk, 2012). In practice, this term is negligible in the present experiments given $|\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}|$ and the resolution of the comparisons. See Appendix A.3 for a further discussion of analyzing the effective sample size for both the class-conditional and prediction-conditional estimates. label-conditional conformal coverage in the special case of class-wise thresholds that only admit prediction sets of cardinality 1. We introduce a straightforward procedure to estimate this quantity next.
3.2.1 Controlling the Class-conditional Accuracy among Selective Classifications with SDM Estimators
In general, the statistical coverage guarantee of marginal split-conformal estimators is not directly informative for selection, since the coverage guarantee is not conditional on the set size. We may instead seek one of various approximately conditional notions of coverage (Romano et al., 2020; Angelopoulos et al., 2021, inter alia); however, there is no guarantee that when we stratify by sets of cardinality 1, coverage will be maintained. However, there is a special case in which label-conditional conformal estimators do provide a meaningful notion of class-conditional coverage for selection. Assuming $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ and $\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}$ are exchangeable, if the conformity score for each label is from a categorical distribution and the resulting thresholding of the class-wise empirical CDFs results in class-wise thresholds that are all greater than $\frac{1}{C}$, then the cardinality 1 sets will, on average, obtain class-conditional coverage, by definition. [^5] Although the formal interpretations are not identical, the evaluation of class-conditional coverage for a single estimate of such cardinality 1 prediction sets in this restricted setting over $\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}$ is numerically equivalent to assessing the class-conditional accuracy, when not considering the sample size error term. Unfortunately, it may be rare to encounter this restricted setting over the full data distributions of real-world tasks. Instead, we will use the $\operatorname{\text{{sdm}}}$ activation to estimate a partitioning of the distribution into a region that approximately fulfills these assumptions.